# Robust Score-Based Quickest Change Detection

### v3

@@ VIII-B Gauss-Bernoulli Restricted Boltzmann Machine
$\nabla_{x}\log q_{\infty}(x)=\sum_{i=0}^{1}\beta_{\infty}^{i}(x)\nabla_{x}\log g_{\infty}^{i}(x):\sum_{i=1}^{m}\beta_{\infty}^{i}(x)=1,\;\beta_{\infty}^{i}(x)\geq 0,\;$ (52)
$\nabla_{x}\log q_{1}(x)=\sum_{i=0}^{1}\beta_{1}^{i}(x)\nabla_{x}\log g_{1}^{i}(x):\sum_{i=1}^{m}\beta_{1}^{i}(x)=1,\;\beta_{1}^{i}(x)\geq 0.\;$ (53)
We use a neural network $\operatorname{Softmax}_{j}\circ f_{\infty}(x)$ to estimate $\beta_{\infty}^{j}(\cdot)$, specifically,
@@
To proceed with Algorithm 1, we need a method of calculating the Hyvärinen scores $\mathcal{S}_{H}(\cdot,Q_{\infty}),\mathcal{S}_{H}(\cdot,Q_{1})$. These Hyvärinen scores are the sum of a gradient log-density, defined in Equations 52 and 53, along with a Laplacian term. To estimate the Laplacian term of a mixture of GBRBM distributions, we use Hutchinson’s Trick ([25, 22]):
@@
$P_{\infty}$ $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})$ $\mathbb{D}_{F}(P_{\infty}\|Q_{1})$
@@
We can see that $\mathbb{D}_{F}(G_{\infty}^{1}\|Q_{\infty})$ is very close to zero, and this gives further evidence that the LFD $Q_{\infty}=G_{\infty}^{1}$. By Theorem VI.3, Assumption VI.1 holds for all $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
Trials R-00, R-01, R-10, and R-11 are robust and use the least-favorable distributions in the calculation of their instantaneous detection scores, while trial N is a nonrobust test and uses non-least-favorable distributions in the calculation of the instantaneous detection scores. As before, the robust drifts are negative before the change point and positive after the change point, but the nonrobust drifts are positive before the change point. Next, we vary the detection threshold $\omega$ and generate plots comparing mean time to false alarm with the expected detection delay:
@@ IX Conclusion
In this work, we proposed the RSCUSUM algorithm, which can detect a change in the distribution of an online high-dimensional data stream in real time. While the SCUSUM algorithm developed in our previous work could accomplish this by knowing only the score function of the pre- and post-change distributions, the new RSCUSUM method can accomplish this with the additional challenge of imprecisely knowing the pre- and post-change distribution score functions.
@@ Appendix A Selection of Appropriate Multiplier
$h(\rho)\overset{\Delta}{=}\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]-1.$
Observe that
$h^{\prime}(\rho)\overset{\Delta}{=}\frac{dh}{d\rho}(\rho)=\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))\exp(z_{\rho}(X_{1}))].$
Note that $h(0)=0$, and $h^{\prime}(0)=\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})<0$ by Assumption VI.1.
@@
$h^{\prime\prime}(\rho)\overset{\Delta}{=}\frac{d^{2}h}{d\rho}(\rho)=\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))^{2}\exp(z_{\rho}(X_{1}))]\geq 0.$
We claim that $h(\rho)$ is strictly convex, namely $h^{\prime\prime}(\rho)>0$ for all $\rho\in[0,\infty)$. Suppose $h^{\prime\prime}(\rho)=0$ for some $\rho\geq 0$, we must have $S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0$ almost surely. This implies that $\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))]=0$ which in turn gives $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})=0$, violating Assumption VI.1. Thus, $h(\rho)$ is strictly convex and $h^{\prime}(\rho)$ is strictly increasing.
@@
    Note that in this case, we must have $P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\geq c\right)>0$, for some $c>0$. Otherwise, we have $P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\leq 0\right)=1$. This implies that $P_{\infty}(z_{\rho}(X_{1})\leq 0)=1$, or equivalently $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]<1$ for all $\rho>0$, and therefore leads to Case 2: $h(\rho)<0$ for all $\rho>0$. Here, $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\neq 1$ since $P_{\infty}(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0)<1$; otherwise $P_{\infty}(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0)=1$, and then $\mathbb{E}_{\infty}[S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})]=\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})=0$, causing the same contradiction to Assumption VI.1.
  • Case 2: If $h(\rho)$ is strictly decreasing in $(0,\infty)$, then any $\rho\in(0,\infty)$ satisfies Inequality (10). As discussed before, in this case, we must have $P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\leq 0\right)=1$. Equivalently, all the increments of the RSCUSUM detection score are non-positive under the pre-change distribution, and $P_{\infty}(Z(n)=0)=1$ for all $n$. Accordingly, $\mathbb{E}_{\infty}[T_{\textit{RSCUSUM}}]=+\infty$. When there occurs change (under measure $Q_{1}$), we also observe that RSCUSUM can get close to detecting the change-point instantaneously as $\rho$ is chosen arbitrarily large. Obviously, this case is of no practical interest.
@@ Appendix B Proofs of Delay and False Alarm Theorems
  1. 1.
    Define
    $Z_{\rho}(0)=0,$
    $Z_{\rho}(n)\overset{\Delta}{=}(Z_{\rho}(n-1)+z_{\rho}(X_{1}))^{+},\;\forall n\geq 1,$
@@
    For convenience, define $\omega^{\prime}\overset{\Delta}{=}\omega\rho$.
  2. 2.
    We next construct a non-negative martingale with mean $1$ under the measure $P_{\infty}$. Define a new instantaneous score function $x\mapsto\tilde{z}_{\rho}(x)$ given by
    $\tilde{z}_{\rho}(x)\overset{\Delta}{=}z_{\rho}(x)+\delta,$
@@
    $\delta\overset{\Delta}{=}-\log\biggr{(}\mathbb{E}_{\infty}\left[\exp(z_{\rho}(X_{1}))\right]\biggr{)}.$
@@
    $\tilde{G}_{n}\overset{\Delta}{=}\exp\biggr{(}\sum_{k=1}^{n}\tilde{z}_{\rho}(X_{k})\biggr{)},\;\forall n\geq 1.$
@@
  3. 3.
    We next examine the new stopping rule
@@
    $\rho(\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1}))$ $=\mathbb{E}_{\infty}[z_{\rho}(X_{1})]=c=\mathbb{E}_{1}[z_{\rho}(X_{1})]\geq\rho\mathbb{D}_{F}(Q_{1}\|Q_{\infty})$
    Recalling Assumption VI.1 and dividing through by $\rho>0$, we know that $0>\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})\geq\mathbb{D}_{F}(Q_{1}\|Q_{\infty})\geq 0$, the last inequality a result of the fact that the Fisher distance is nonnegative. Thus, the inequality of Equation (61) is strict. Continuing from Equation 60:
@@
    where the strictness of the final inequality follows from the strictness of the inequality in Equation (61). Taking the log, we have that $\delta<-\rho(\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1}))$, so
    $\mathbb{E}_{\infty}[\tilde{z}_{\rho}(X_{i})]=\mathbb{E}_{\infty}[z_{\rho}(X_{i})]+\delta<\rho(\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1}))-\rho(\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1}))=0$
@@
  5. 5.
    Define a sequence of stopping times:
    $\eta_{0}$ $=0,$
    $\eta_{1}$ $=\inf\left\{t:\sum_{i=1}^{t}\tilde{z}_{\rho}(X_{i})<0\right\},$
    $\eta_{k+1}$ $=\inf\left\{t>\eta_{k}:\sum_{i=\eta_{k}+1}^{t}\tilde{z}_{\rho}(X_{i})<0\right\},\;\text{for}\;k\geq 1.$
@@
    $M\overset{\Delta}{=}\inf\biggl{\{}k\geq 0:\eta_{k}<\infty\;\text{and}\;\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\biggr{\}}.$ (62)
@@
  6. 6.
    For the probability $P_{\infty}(M>k)$, note that since the event $\{M<k\}$ is $\mathcal{F}_{\eta_{k}}-$measurable,
    $\begin{split}P_{\infty}(M>k)&=\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}]=\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}\mathbb{I}_{\{M\geq k\}}]=\mathbb{E}_{\infty}\left[\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}\mathbb{I}_{\{M\geq k\}}|\mathcal{F}_{\eta_{k}}]\right]\\ &=\mathbb{E}_{\infty}[P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})\mathbb{I}_{\{M\geq k\}}].\end{split}$ (63)
@@
    $P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})=1-P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right).$ (64)
@@
    $P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right)=P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right).$
@@
    $\begin{split}P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right)&=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})>\omega^{\prime}\right)\\ &=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\exp\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\right)>e^{\omega^{\prime}}\right)\\ &=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\tilde{G}_{n}>e^{\omega^{\prime}}\right).\end{split}$
@@
    $\begin{split}P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})&=1-P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right)\\ &=1-P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right)\\ &=1-\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\tilde{G}_{n}>e^{\omega^{\prime}}\right)\geq 1-e^{\omega^{\prime}}.\end{split}$ (66)
  7. 7.
    Finally, from (63), we have
    $\begin{split}P_{\infty}(M>k)&=\mathbb{E}_{\infty}[P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})\mathbb{I}_{\{M\geq k\}}]\\ &\geq(1-e^{\omega^{\prime}})P_{\infty}(M>k-1)\\ &\geq(1-e^{\omega^{\prime}})^{2}P_{\infty}(M>k-2)\\ &\geq\dots\geq(1-e^{\omega^{\prime}})^{k}\end{split}$ (67)
@@
$T_{\texttt{RSCUSUM}}^{\prime}\overset{\Delta}{=}\inf\{n\geq 1:Z^{\prime}(n)\geq\omega\}.$
@@
$R_{\omega}\overset{\Delta}{=}Z^{\prime}(T_{\texttt{RSCUSUM}}^{\prime})-\omega.$
$R_{\omega}$ is the excess of the random walk over a stopping threshold $\omega$ at the stopping time $T_{\texttt{RSCUSUM}}^{\prime}$. Suppose the change-point $\nu=1$, then $X_{1},X_{2},\ldots,$ are i.i.d. following the distribution $Q_{1}$. Let $\mu$ and $\sigma^{2}$ respectively denote the mean $\mathbb{E}_{1}[z(X_{1})]$ and the variance $\text{Var}_{1}[z(X_{1})]$. Note that
@@
Under the mild regularity conditions given by [16], which we assume in Assumption IV.4:
$\mathbb{E}_{1}[\mathcal{S}_{\texttt{H}}(X_{1},Q_{\infty})^{2}]<\infty,\;\text{and}$
$\mathbb{E}_{1}[\mathcal{S}_{\texttt{H}}(X_{1},Q_{1})^{2}]<\infty.$
Therefore, by [27] Theorem 1,

### v4

@@ VIII-B Gauss-Bernoulli Restricted Boltzmann Machine
$\nabla_{x}\log q_{\infty}(x)=\sum_{i=0}^{1}\beta_{\infty}^{i}(x)\nabla_{x}\log g_{\infty}^{i}(x),$ (55)
where $\sum_{i=1}^{m}\beta_{\infty}^{i}(x)=1,\;\beta_{\infty}^{i}(x)\geq 0$, and
$\nabla_{x}\log q_{1}(x)=\sum_{i=0}^{1}\beta_{1}^{i}(x)\nabla_{x}\log g_{1}^{i}(x),$ (56)
where $\sum_{i=1}^{m}\beta_{1}^{i}(x)=1,\;\beta_{1}^{i}(x)\geq 0$. We use a neural network $\operatorname{Softmax}_{j}\circ f_{\infty}(x)$ to estimate $\beta_{\infty}^{j}(\cdot)$, specifically,
@@
To proceed with Algorithm 1, we need a method of calculating the Hyvärinen scores $\mathcal{S}_{\texttt{{H}}}(\cdot,Q_{\infty}),\mathcal{S}_{\texttt{{H}}}(\cdot,Q_{1})$. These Hyvärinen scores are the sum of a gradient log-density, defined in Equations 55 and 56, along with a Laplacian term. To estimate the Laplacian term of a mixture of GBRBM distributions, we use Hutchinson’s Trick ([25, 22]):
@@
$P_{\infty}$ $\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})$ $\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})$
@@
We can see that $\mathbb{D}_{\texttt{{F}}}(G_{\infty}^{1}\|Q_{\infty})$ is very close to zero, and this gives further evidence that the LFD $Q_{\infty}=G_{\infty}^{1}$. By Theorem VI.3, Assumption VI.1 holds for all $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
Trials R-00, R-01, R-10, and R-11 are robust and use the least-favorable distributions in the calculation of their instantaneous detection scores, while trial N is a nonrobust test and uses non-least-favorable distributions in the calculation of the instantaneous detection scores. As before, the robust drifts are negative before the change point and positive after the change point, but the nonrobust drifts are positive before the change point.
Next, we vary the detection threshold $\omega$ and generate plots comparing mean time to false alarm with the expected detection delay.
@@ IX Conclusion
In this work, we proposed the RSCUSUM algorithm, which can detect a change in the distribution of an online high-dimensional data stream in real time. While the SCUSUM algorithm could accomplish this by knowing only the score function of the pre- and post-change distributions, the new RSCUSUM method can accomplish this with the additional challenge of imprecisely knowing the pre- and post-change distribution score functions.
@@ Appendix A Selection of Appropriate Multiplier
$h(\rho)=\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]-1.$
where $z_{\rho}(x)$ is defined as in (27). Observe that
$h^{\prime}(\rho)=\frac{dh}{d\rho}(\rho)\\ =\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))\exp(z_{\rho}(X_{1}))].$
Note that $h(0)=0$, and $h^{\prime}(0)=\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})<0$ by Assumption VI.1.
@@
$h^{\prime\prime}(\rho)=\frac{d^{2}h}{d\rho}(\rho)\\ =\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))^{2}\exp(z_{\rho}(X_{1}))]\geq 0.$
We claim that $h(\rho)$ is strictly convex, namely $h^{\prime\prime}(\rho)>0$ for all $\rho\in[0,\infty)$. Suppose $h^{\prime\prime}(\rho)=0$ for some $\rho\geq 0$, we must have $S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0$ almost surely. This implies that
$\mathbb{E}_{\infty}[(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1}))]=0$
which in turn gives
$\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})=0,$
violating Assumption VI.1. Thus, $h(\rho)$ is strictly convex and $h^{\prime}(\rho)$ is strictly increasing.
@@
Note that in this case, we must have
$P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\geq c\right)>0$
for some $c>0$. Otherwise, we have
$P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\leq 0\right)=1.$
This implies that $P_{\infty}(z_{\rho}(X_{1})\leq 0)=1$, or equivalently $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]<1$ for all $\rho>0$, and therefore leads to Case 2: $h(\rho)<0$ for all $\rho>0$. Here, $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\neq 1$ since
$P_{\infty}(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0)<1;$
otherwise
$P_{\infty}(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})=0)=1,$
and then
$\mathbb{E}_{\infty}[S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})]\\ =\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})=0,$
causing the same contradiction to Assumption VI.1.
Case 2: If $h(\rho)$ is strictly decreasing in $(0,\infty)$, then any $\rho\in(0,\infty)$ satisfies Inequality (10). As discussed before, in this case, we must have
$P_{\infty}\left(S_{\texttt{H}}(X_{1},Q_{\infty})-S_{\texttt{H}}(X_{1},Q_{1})\leq 0\right)=1.$
Equivalently, all the increments of the RSCUSUM detection score are non-positive under the pre-change distribution, and $P_{\infty}(Z(n)=0)=1$ for all $n$. Accordingly, $\mathbb{E}_{\infty}[T_{\textit{RSCUSUM}}]=+\infty$. When there occurs change (under measure $P_{1}$), we also observe that RSCUSUM can get close to detecting the change-point instantaneously as $\rho$ is chosen arbitrarily large. Obviously, this case is of no practical interest.
@@ Appendix B Proofs of Delay and False Alarm Theorems
Part 1: Define
$Z_{\rho}(0)=0,\quad Z_{\rho}(n)=(Z_{\rho}(n-1)+z_{\rho}(X_{1}))^{+},\;\forall n\geq 1,$
@@
For convenience, define $\omega^{\prime}=\omega\rho$.
Part 2: We next construct a non-negative martingale with mean $1$ under the measure $P_{\infty}$. Define a new instantaneous score function $x\mapsto\tilde{z}_{\rho}(x)$ given by
$\tilde{z}_{\rho}(x)=z_{\rho}(x)+\delta,$
@@
$\delta=-\log\biggr{(}\mathbb{E}_{\infty}\left[\exp(z_{\rho}(X_{1}))\right]\biggr{)}.$
@@
$\tilde{G}_{n}=\exp\biggr{(}\sum_{k=1}^{n}\tilde{z}_{\rho}(X_{k})\biggr{)},\;\forall n\geq 1.$
@@
Part 3: We next examine the new stopping rule
@@
$\rho(\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1}))=\mathbb{E}_{\infty}[z_{\rho}(X_{1})]\\ =c=\mathbb{E}_{1}[z_{\rho}(X_{1})]\geq\rho\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty})$
Recalling Assumption VI.1 and dividing through by $\rho>0$, we know that $0>\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})\geq\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty})\geq 0$, the last inequality a result of the fact that the Fisher distance is nonnegative. Thus, the inequality of Equation (64) is strict. Continuing from Equation 63:
@@
where the strictness of the final inequality follows from the strictness of the inequality in Equation (64). Taking the log, we have that $\delta<-\rho(\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1}))$, so
$\mathbb{E}_{\infty}[\tilde{z}_{\rho}(X_{i})]$ $=\mathbb{E}_{\infty}[z_{\rho}(X_{i})]+\delta<\rho(\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1}))-\rho(\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1}))=0$
@@
Part 5: Define a sequence of stopping times:
$\eta_{0}=0,\quad\quad\eta_{1}=\inf\left\{t:\sum_{i=1}^{t}\tilde{z}_{\rho}(X_{i})<0\right\},\quad\quad\eta_{k+1}=\inf\left\{t>\eta_{k}:\sum_{i=\eta_{k}+1}^{t}\tilde{z}_{\rho}(X_{i})<0\right\},\;\text{for}\;k\geq 1.$
@@
$M=\inf\biggl{\{}k\geq 0:\eta_{k}<\infty\\ \;\text{and}\;\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\biggr{\}}.$ (65)
@@
Part 6: For the probability $P_{\infty}(M>k)$, note that since the event $\{M<k\}$ is $\mathcal{F}_{\eta_{k}}-$measurable,
$P_{\infty}(M>k)=\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}]=\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}\mathbb{I}_{\{M\geq k\}}]$ $=\mathbb{E}_{\infty}\left[\mathbb{E}_{\infty}[\mathbb{I}_{\{M>k\}}\mathbb{I}_{\{M\geq k\}}|\mathcal{F}_{\eta_{k}}]\right]$
$=\mathbb{E}_{\infty}[P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})\mathbb{I}_{\{M\geq k\}}].$ (66)
@@
$P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})=\\ 1-P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right).$ (67)
@@
$P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right)\\ =P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right).$
@@
$P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right)$ $=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})>\omega^{\prime}\right)$
$=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\exp\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\right)>e^{\omega^{\prime}}\right)$
$=\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\tilde{G}_{n}>e^{\omega^{\prime}}\right).$
@@
$P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})$ $=1-P_{\infty}\left(\sum_{i=\eta_{k}+1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n>\eta_{k}\mid\mathcal{F}_{\eta_{k}}\right)$
$=1-P_{\infty}\left(\sum_{i=1}^{n}\tilde{z}_{\rho}(X_{i})\geq\omega^{\prime}\;\text{for some}\;n\right)$
$=1-\lim_{m\rightarrow\infty}P_{\infty}\left(\max_{n:n\leq m}\tilde{G}_{n}>e^{\omega^{\prime}}\right)$
$\geq 1-e^{-\omega^{\prime}}.$ (69)
Part 7: Finally, from (B-A), we have
$\begin{split}P_{\infty}(M>k)&=\mathbb{E}_{\infty}[P_{\infty}(M\geq k+1\mid\mathcal{F}_{\eta_{k}})\mathbb{I}_{\{M\geq k\}}]\\ &\geq(1-e^{-\omega^{\prime}})P_{\infty}(M>k-1)\\ &\geq(1-e^{-\omega^{\prime}})^{2}P_{\infty}(M>k-2)\\ &\geq\dots\geq(1-e^{-\omega^{\prime}})^{k}\end{split}$ (70)
@@
$T_{\texttt{{RSCUSUM}}}^{\prime}=\inf\{n\geq 1:Z^{\prime}(n)\geq\omega\}.$
@@
$R_{\omega}=Z^{\prime}(T_{\texttt{{RSCUSUM}}}^{\prime})-\omega.$
$R_{\omega}$ is the excess of the random walk over a stopping threshold $\omega$ at the stopping time $T_{\texttt{{RSCUSUM}}}^{\prime}$. Suppose the change-point $\nu=1$, then $X_{1},X_{2},\ldots,$ are i.i.d. following the distribution $P_{1}$. Let $\mu$ and $\sigma^{2}$ respectively denote the mean $\mathbb{E}_{1}[z(X_{1})]$ and the variance $\text{Var}_{1}[z(X_{1})]$. Note that
@@
Under the mild assumptions that:
$\mathbb{E}_{1}[\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{\infty})^{2}]<\infty,\;\quad\mathbb{E}_{1}[\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{1})^{2}]<\infty,$
then by [27, Theorem 1],